OpenStudy (anonymous):

Given triangle QRS is congruent to triangle TUV, QS=3v+2 and TV=7v-6, find the length of QS and TV. Please help I don't know where to start. @Yuba

OpenStudy (anonymous):

Im pretty sure you get an image...right?o.o

OpenStudy (anonymous):

No, I don't.

OpenStudy (anonymous):

But I must show my work. For it is a "WorkPad" problem.

OpenStudy (anonymous):

Alright, since we know congruent = same shape and size, that means: Q = T R = U and S = V Now we know that QS = TV so, 3v + 2 = 7v - 6

OpenStudy (anonymous):

Is that the answer? Or do I have to solve something?

OpenStudy (anonymous):

Do you know how to solve that equation?

OpenStudy (anonymous):

Do you subtract 3v from both sides first?

OpenStudy (anonymous):

Yes you do :) So then you have: 2 = 4v - 6 Then you add 6 to both sides so: 8 = 4v

OpenStudy (anonymous):

So 2=v?

OpenStudy (anonymous):

Yup, v = 2 :3 now find the length of both sides TV and QS Plug in the variable 3(2) + 2 = 7(2) - 6 6 + 2 = 14 - 6 8 = 8 :)

OpenStudy (anonymous):

And that's the final answer?

OpenStudy (anonymous):

Yup :3 Both lengths are 8

OpenStudy (anonymous):

Wow, thank you so much!

OpenStudy (anonymous):

I hate math >-> Havent done it in over 6 months xD You're welcome though <3 :3

OpenStudy (anonymous):

But yet you still managed to help me out. :P